OpenStudy (anonymous):

Can anyone help me with Lagrange Multipliers? find the point in the first quadrant on the curve xy^4=31 that is closest to the origin.

OpenStudy (anonymous):

use x^2+y^2+z^2 -a^2 for your distance function grad (x^2+y^2+z^2 -a^2) = lambda *grad(xy^4-31) find the lambdas that give positive x and y

OpenStudy (anonymous):

where is the z^2-a^2 coming from?

OpenStudy (anonymous):

it's a general form for 3 space. you'll use it for any function g(x,y,z). here you notice that it won't matter because there's no k coordinate of the gradient of your curve, so 2z=0

OpenStudy (anonymous):

alright so i did that and i got the 3 equations but i get stuck. i was able to do a problem simmilar but there was not both x and y in a partial.

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

counting 2z=0 ?

OpenStudy (anonymous):

umm no ill type out what i have one moment

OpenStudy (anonymous):

2x = lambda(y^4) 2y= lambda(4xy^3)

OpenStudy (anonymous):

\[2x=\lambda(xy^4-31)\] \[2y=\lambda(y^4-31)\] \[xy^4=31\]

OpenStudy (anonymous):

whoops your right i dident mean to add the 31's in any of those

OpenStudy (anonymous):

what's the gradient of xy^4-31

OpenStudy (anonymous):

<y^4,4xy^3> i really messed up typing those lol sorry

OpenStudy (anonymous):

+ 0 k yes, right.

OpenStudy (anonymous):

so using those equations should i love for lambda first?or x and y?

OpenStudy (anonymous):

solve the second for y^2 and sub.s into the first.

OpenStudy (anonymous):

so y^2=4((lambda)^2)(x^6)?

OpenStudy (anonymous):

you don't have to do it that way, that's just what I'd do..

OpenStudy (anonymous):

y^2 = 1/(lambda*2x)

OpenStudy (anonymous):

ehhh i dont understand what im trying to get to. in the end i need lambda, x, and y right?

OpenStudy (anonymous):

lambda, really...

OpenStudy (anonymous):

if you find what values lambda must have it tells you how x and y must be related for your constraint to hold.

OpenStudy (anonymous):

alright i found x=lambda(y^4)/2 and y=\[\sqrt[8]{62/\lambda}\]

OpenStudy (anonymous):

but for some reason i feel like im making this harder then it needs to be

OpenStudy (anonymous):

looks like you're going to have to sub.s back into the original equation to find lambda for this problem..

OpenStudy (anonymous):

ok i thought so i started to do that and i got \[(x^2\lambda2x+1)/\lambda2x\]